OF a mixture containing 7 quarts, 5 quarts are alcohol. how many quarts of water must be added to produce a mixure containing 25%alcohol?
Found 2 solutions by stanbon, nerdybill:
Answer by stanbon(75887)

OF a mixture containing 7 quarts, 5 quarts are alcohol. how many quarts of water must be added to produce a mixure containing 25%alcohol?
-----------------------------------------
Equation:
alcohol + alcohol = alcohol
5 + 0*x = 0.25(7+x)
5 = (7/4) + (x/4)
20 = 7 + x
x = 13 quarts (amount of water needed in the mixture)
===========================
Cheers,
Stan H.

Answer by nerdybill(7384)

OF a mixture containing 7 quarts, 5 quarts are alcohol. how many quarts of water must be added to produce a mixure containing 25%alcohol?
.
Let x = amount of water added
then
5 = .25(7+x)
20 = 7+x
13 quarts = x